# Nonuniqueness of capped black holes: large and small bubbles

### v1

@@ I Introduction
In this paper, using the Ehlers-Harrison transformations, previously applied to the charged rating black ring with all independent conserved charges Suzuki:2024vzq , we wish to generalize the three-parameter family of the capped black hole solution, derived in our previous works Suzuki:2023nqf ; Suzuki:2024phv , to a four-parameter family. The Ehlers and Harrison transformations add an angular momentum and an electric charge to a five-dimensional vacuum solution, respectively, preserving asymptotic flatness. In our previous work, we developed the construction of a vacuum black ring solution possessing a Dirac-Misner string singularity through the Ehlers transformation, demonstrating that a doubly rotating black ring (Pomeransky-Sen’kov solution) can be derived once the Dirac-Misner string singularity is removed. Subsequently, in Ref. Suzuki:2024vzq , we applied the Harrison transformation to this vacuum solution, generating a corresponding charged solution within five-dimensional minimal supergravity. By setting appropriate boundary conditions on the parameters, we successfully eliminated all singularities—including those related to curvature, conical features, and Dirac-Misner strings—across the rotational axes and the horizon. Using this methodology, we can construct a capped black hole, distinguishing it by replacing the $S^{1}\times S^{2}$ topology with $S^{3}$. We demonstrate that the new solution manifests discrete non-uniqueness in terms of conserved charges, presenting two distinct branches that share the same mass, angular momenta, and electric charges as the Cvetič-Youm black hole. These branches are distinguished by the size of their bubbles adjacent to the horizon and are accordingly termed small bubble and large bubble branches. Our analysis indicates that within one parameter region, the large bubble branch exhibits greater entropy than both the small bubble branch and the Cvetič-Youm black hole, making it thermodynamically more favored. Conversely, in another parameter region, the Cvetič-Youm black hole displays higher entropy than both the large and small bubble branches.
The rest of the paper is organized as follows. In section II, we briefly review the derivation of the solution for the metric and gauge potential, and present their expression in the C-metric coordinates. In section III, we clarify the boundary condition for the absence of curvature, conical, orbifold, Dirac-Misner string singularities and closed timelike curves. In section IV, we investigate the phase of the physical solution. Finally, we summarize the results in section V.
@@ II.1 Setup
Assuming that there is one timelike Killing vector $\xi_{0}=\partial/\partial t$ and one spacelike axial Killing vector $\xi_{1}=\partial/\partial\psi$, this theory simplifies to the $G_{2(2)}/SL(2,{\mathbb{R}})\times SL(2,{\mathbb{R}})$ non-linear sigma models coupled with three-dimensional gravity Mizoguchi:1998wv ; Mizoguchi:1999fu . Following the discussion in Ref. Tomizawa:2009ua , from Eqs. (2) and (3), one can define the gravitational potential $\lambda_{ab}$, electric potential $\psi_{a}$, magnetic potential $\mu$ and twist potentials $\omega_{a}$ by
$\lambda_{ab}$ $=$ $(\xi_{a}|\xi_{b}),$ (4)
$d\psi_{a}$ $=$ $-i_{\xi_{a}}F,$ (5)
$d\mu$ $=$ $\frac{1}{\sqrt{3}}\star(\xi_{0}\wedge\xi_{1}\wedge F)-\epsilon^{ab}\psi_{a}d\psi_{b}\,,$ (6)
$d\omega_{a}$ $=$ $\star(\xi_{0}\wedge\xi_{1}\wedge d\xi_{a})+\psi_{a}(3d\mu+\epsilon^{bc}\psi_{b}d\psi_{c})\,,$ (7)
where $\epsilon^{01}=-\epsilon^{10}=1$. Then, the action (1) reduces to the nonlinear sigma model for the eight scalar functions $\{\lambda_{ab},\omega_{a},\psi_{a},\mu\}$ coupled with three-dimensional gravity.
Furthermore, if we assume the existence of a third spacelike axial Killing vector $\xi_{2}=\partial/\partial\phi$ commuting with the other two, the metric can be expressed in the Weyl-Papapetrou form
$ds^{2}$ $=$ $\lambda_{ab}(dx^{a}+a^{a}{}_{\phi}{}d\phi)(dx^{b}+a{}^{b}{}_{\phi}d\phi)+\tau^{-1}\rho^{2}d\phi^{2}+\tau^{-1}e^{2\sigma}(d\rho^{2}+dz^{2})\,,$ (8)
and the gauge potential is written,
$A=\sqrt{3}\psi_{a}dx^{a}+A_{\phi}d\phi\,,$ (9)
where the coordinates $x^{a}=(t,\psi)$ ($a=0,1$) denote the Killing coordinates, and thus all functions $\lambda_{ab}$, $\tau:=-{\rm det}(\lambda_{ab})$, $a^{a}$, $\sigma$, and $(\psi_{a},A_{\phi})$ are independent of $\phi$ and $x^{a}$. Note that the coordinates $(\rho,z)$ that span a two-dimensional base space, $\Sigma=\{(\rho,z)|\rho\geq 0,\ -\infty<z<\infty\}$, are globally well-defined, harmonic, and mutually conjugate on $\Sigma$. The metric functions $a^{a}{}_{\phi}\ (a=0,1)$ and the component $A_{\phi}$ of the gauge potential are determined by the eight scalar functions $\{\lambda_{ab},\omega_{a},\psi_{a},\mu\}$ from Eqs. (6) and (7), and the function $\sigma$ is determined from these scalar functions up to a constant factor. Therefore, the theory reduces to the nonlinear sigma model of $\{\lambda_{ab},\omega_{a},\psi_{a},\mu\}$ decoupled from the gravity. Because the theory is invariant under the $G_{2(2)}$-transformation, one can produce a new solution from a known solution by applying a certain transformation in the $G_{2(2)}$-group. Here, we focus on two different $G_{2(2)}$-transformations.
One is a transformation in the gravitational sector that reduces to $SL(3,{\mathbb{R}})/SL(2,{\mathbb{R}})$ non-linear sigma model for the vacuum solution $\Phi^{A}=\{\lambda_{ab},\omega_{a},\psi_{a}=0,\mu=0\}$, which is invariant under the $SL(3,{\mathbb{R}})$ transformation Maison:1979kx . Giusto and Saxena found that a $SL(2,{\mathbb{R}})$ subgroup in $SL(3,{\mathbb{R}})$, or the Ehlers transformation, can be utilized to add angular momenta to five-dimensional asymptotically-flat, vacuum solutions Giusto:2007fx . With this transformation, one can obtain the new metric describing the rotating solution to the vacuum Einstein equations with the same Killing isometries. For example, in Ref. Giusto:2007fx , applying this transformation to the five-dimensional static black hole solution (the five-dimensional Schwarzschild-Tangherlini solution Tangherlini:1963bw ) one can reproduce the five-dimensional rotating black hole solution (the five-dimensional Myers-Perry solution Myers:1986un ). However, when applied to vacuum black rings, such as the static black ring and the singly rotating black ring, a Dirac-Misner string singularity inevitably appears on the disk inside the ring, similar to the Harrison transformation studied in Ref. Bouchareb:2007ax . This issue is solved by using the seed with a Dirac-Misner string singularity. In our prior study Suzuki:2024eoz , applying the Ehlers transformation to the vacuum solution with a Dirac-Misner string singularity constructed in Ref. Chen:2008fa , we can successfully reproduce the known doubly-rotating vacuum black ring solution (Pomeransky-Senkov solution) Pomeransky:2006bd .
The other is the so-called Harrison transformation formulated by Bouchareb et. al. in Ref. Bouchareb:2007ax . This transforms a five-dimensional vacuum solution $\Phi^{A}=\{\lambda_{ab},\omega_{a},\psi_{a}=0,\mu=0\}$ into a charged solution $\Phi^{\prime A}=\{\lambda^{\prime}_{ab},\omega^{\prime}_{a},\psi^{\prime}_{a},\mu^{\prime}\}$ in the five-dimensional minimal supergravity while preserving asymptotic flatness. In particular, this transformation adds the electric and dipole charges to a vacuum black ring solution while keeping the asymptotic flatness and Killing isometries. However, as mentioned in Ref. Bouchareb:2007ax , when one performs the Harrison transformation for the regular vacuum black ring such as the Pomeransky-Sen’kov solution, a Dirac-Misner string singularity inevitably appears on the disk inside the ring, though the transformation can generate the regular Cvetič-Youm charged black hole for the vacuum black hole such as the Myers-Perry solution. Instead, performing the Harrison transformation to the vacuum solution with a Dirac-Misner string singularity, one can obtain regular charged black holes such as the capped black hole Suzuki:2023nqf ; Suzuki:2024phv and charged black rings Suzuki:2024coe ; Suzuki:2024vzq .
II.2 Charged solution
@@ III Boundary conditions
In order that the solution (10) and (II.2) describes a capped black hole of physical interest, we must impose the appropriate boundary conditions at
@@ III.1 Asymptotic infinity ($x\to y\to-1$)
From the asymptotic behavior of the gauge potential (II.2), one can determine the electric charge
@@ III.2 Rotational Axes ($y=-1,x=\pm 1$)
    To specifically ensure a topology of $S^{3}$, we must set $n=\pm 1$. This work focuses on $n=1$.
@@
    The condition for the absence of conical singularities: Since on this axis, the Killing vector $\partial_{\phi_{+}}:=\partial_{\phi}+\partial_{\psi}$ vanishes under the two conditions previously mentioned, the inner axis does not exhibit a conical singularity if
@@ III.3 The center $(x,y)=(1,-1)$
where $t^{\prime}:=\sqrt{(1-\gamma)(1+\nu)}f_{1}f_{2}t/(c^{2}(1+\nu)f_{2}^{2}-s^{2}(1-\gamma)f_{1}^{2})$ and $(\psi_{+},\phi_{+})=(\psi-\phi,\phi)$. Since the inequalities (25), (51) and (53) guarantees ${\mathcal{C}}_{1}>0$, the spacetime metric becomes the Minkowski metric. Therefore, the point $(x,y)=(1,-1)$ turns out to be a regular center of the Minkowski spacetime.
@@ III.4 Horizon
The spatial cross section of the horizon has topology $S^{3}$ under the boundary condition (40) with $n=1$. The area of the horizon cross section at $y=-1/\nu$ is given in terms of the surface gravity
@@
where $\omega_{\psi}^{+}=\omega_{\psi}-\omega_{\phi}$ , $\omega_{\psi}^{-}=\omega_{\psi}$ , and $(\psi_{\pm},\phi_{\pm})$ are defined in Eq. (47). $R_{\pm}:=\sqrt{g_{\psi\psi}}|_{(x,y)=(\pm 1,-1/\nu)}$ are given by
@@ III.5 Absence of curvature singularities and CTCs
Figure 1: Profile of $H(x,y)$ for $(x,y)\in[-1,1]\times[-1/\nu,-1]$ and $(\sigma,\tilde{\nu},\tilde{\gamma})=(0.8,1.2,0.239)$.
@@
Figure 2: Profile of ${\cal D}(x,y)$ for $(x,y)\in[-1,1]\times[-1/\nu,-1]$ and $(\sigma,\tilde{\nu},\tilde{\gamma})=(0.8,1.2,0.239)$.
@@ III.6 Finding parameters for capped black holes
We are now aiming to find a capped black hole solution that does not have any curvature singularities, orbifold singularities, conical singularities, or Dirac-Misner string singularities both on and outside the horizon, additionally, they are free from closed timelike curves. This can be achieved if we can find a set of parameters $(\ell,\nu,\gamma,a,b,\beta,\alpha)$ that meets three constraints (39), (40) and (41). As outlined in our previous works Suzuki:2023nqf ; Suzuki:2024phv , such black holes are distinguished by having a compact inner rotational axis with a two-dimensional disk topology, referred to as a “bubble,” which lies adjacent to the horizon. The existence of this extra bubble implies that the intersection of the DOC and timeslice $\Sigma$ has the nontrivial topology of ${\rm DOC}\cap\Sigma=[{\mathbb{R}}^{4}\#{\mathbb{C}}{\mathbb{P}}^{2}]\setminus{\mathbb{B}}^{4}$, differing it from the Cvetič-Youm black hole with the trivial topology of ${\rm DOC}\cap\Sigma={\mathbb{R}}^{4}\setminus{\mathbb{B}}^{4}$ Cvetic:1996xz .
@@
$\sigma^{3}=\frac{\tilde{b}\tilde{\nu}\left(\tilde{b}\tilde{\gamma}\tilde{\nu}+\tilde{b}^{2}\left(\tilde{\gamma}^{2}-2\right)+\left(\tilde{b}^{2}-1\right)\tilde{\nu}^{2}+2\right)}{\tilde{\gamma}\left(\tilde{b}\left(\tilde{b}^{2}-2\right)\tilde{\gamma}\tilde{\nu}+\left(1-2\tilde{b}^{2}\right)\tilde{\nu}^{2}+2\left(\tilde{b}^{2}-1\right)\right)}\frac{f_{2}\tilde{g}_{1}}{f_{1}g_{12}}.$ (87)
@@
Therefore, to obtain a physical solution, one must solve Eqs. (84) and (87) under the conditions (25) and (52) for the parameters $(\tilde{\gamma},\tilde{\nu},\tilde{b},\sigma,\beta)$ within the specified range
@@
Since Eqs. (84) and (87) yield quartic and sixth-order equations of $\beta$, respectively, finding the analytic solution with respect to $\beta$ is not easy. Therefore, we resort to exploring the solutions numerically. Since $\beta$ and $\tilde{b}$ are represented as the functions of $\sigma,\tilde{\nu}$ and $\tilde{\gamma}$ through Eqs. (84) and (87), the physical solution is uniquely determined by a given $(\sigma,\tilde{\nu},\tilde{\gamma})$ in the range (89). In Fig. 3, we illustrate all possible solution for $\sigma=0.8,\ \tilde{\nu}=1.2$ and $0<|\tilde{\gamma}|<\sqrt{2-\tilde{\nu}^{2}}$. Moreover, we find that the metric and gauge potential is invariant under the transformation $\tilde{\gamma}\to-\tilde{\gamma}$, despite $\beta,\tilde{b}$ themselves not being invariant [^1] The metric function $H(x,y),F(x,y),J(x,y)$ are modified by the transformation $\tilde{\gamma}\to-\tilde{\gamma}$ but only upto a common factor $H(x,y)\to CH(x,y),\ F(x,y)\to CF(x,y),\ J(x,y)\to CJ(x,y)$. . As in the black ring case Suzuki:2024vzq , this might imply the possible extension from the $\tilde{\gamma}=0$ phase into the parameter region $\gamma\ (=1-\tilde{\gamma}^{2})>1$ where $\tilde{\gamma}$ becomes pure imaginary and $\beta,\tilde{b}$ become complex. In this paper, however, we restrict ourselves for the phases with all the parameters real.
@@
Figure 3: All solutions for $\sigma=0.8$ and $\tilde{\nu}=1.2$ shown in the $(\tilde{\gamma},\beta)$ and $(\tilde{\gamma},\tilde{b})$ plains. $\beta$ diverges at $\tilde{\gamma}_{\infty}=-0.195...$ and the spacetime seems to be singular there. But the phases are smoothly continued at $\tilde{\gamma}=\tilde{\gamma}_{\infty}$ if we use the other parameter $\beta^{\prime}=\beta^{-1}$ around there. Despite $\beta=\beta(\tilde{\gamma})$ and $\tilde{b}=\tilde{b}(\tilde{\gamma})$ do not exhibit the mirror symmetry with respect to $\tilde{\gamma}$, the metric and gauge functions coincide for $\tilde{\gamma}=C>0$ (blue curve) and $\tilde{\gamma}=-C$ (red curve).

### v2

@@ I Introduction
In this paper, using the Ehlers-Harrison transformations, previously applied to the charged rating black ring with all independent conserved charges Suzuki:2024vzq , we wish to generalize the three-parameter family of the capped black hole solution, derived in our previous works Suzuki:2023nqf ; Suzuki:2024phv , to a four-parameter family. The Ehlers and Harrison transformations add an angular momentum and an electric charge to a five-dimensional vacuum solution, respectively, preserving asymptotic flatness. In our previous work, we developed the construction of a vacuum black ring solution possessing a Dirac-Misner string singularity through the Ehlers transformation, demonstrating that a doubly rotating black ring (Pomeransky-Sen’kov solution Pomeransky:2006bd ) can be derived once the Dirac-Misner string singularity is removed. Subsequently, in Ref. Suzuki:2024vzq , we applied the Harrison transformation to this vacuum solution, generating a corresponding charged solution within five-dimensional minimal supergravity. By setting appropriate boundary conditions on the parameters, we successfully eliminated all singularities—including those related to curvature, conical features, and Dirac-Misner strings—across the rotational axes and the horizon. Using this methodology, we can construct a capped black hole, distinguishing it by replacing the $S^{1}\times S^{2}$ topology with $S^{3}$. We demonstrate that the new solution manifests discrete non-uniqueness in terms of conserved charges, presenting two distinct branches that share the same mass, angular momenta, and electric charges as the Cvetič-Youm black hole. These branches are distinguished by the size of their bubbles adjacent to the horizon and are accordingly termed small bubble and large bubble branches. Our analysis indicates that within one parameter region, the large bubble branch exhibits greater entropy than both the small bubble branch and the Cvetič-Youm black hole, making it thermodynamically more favored. Conversely, in another parameter region, the Cvetič-Youm black hole displays higher entropy than both the large and small bubble branches.
The rest of the paper is organized as follows. In section II, we present the metric and gauge potential for the new non-BPS solution of the capped black hole in the C-metric coordinates. In section III, we impose the boundary conditions for the absence of curvature, conical, orbifold, Dirac-Misner string singularities and closed timelike curves on the solution’s parameters. In section IV, we investigate the phase of the physical solution. Finally, we summarize the results in section V.
@@ II Solution generation
Assuming that there is one timelike Killing vector $\xi_{0}=\partial/\partial t$ and one spacelike axial Killing vector $\xi_{1}=\partial/\partial\psi$, this theory simplifies to the $G_{2(2)}/SL(2,{\mathbb{R}})\times SL(2,{\mathbb{R}})$ non-linear sigma models coupled with three-dimensional gravity Mizoguchi:1998wv ; Mizoguchi:1999fu . By further assuming the existence of a third spacelike axial Killing vector $\xi_{2}=\partial/\partial\phi$ commuting with the other two, the theory reduces to the $G_{2(2)}/[SL(2,\mathbb{R})\times SL(2,\mathbb{R})]$ nonlinear sigma model. Because the theory is invariant under the $G_{2(2)}$-transformation, one can produce a new solution from a known solution by applying a certain transformation in the $G_{2(2)}$-group. Below, we present the solution derived through two different $G_{2(2)}$-transformation: the Ehlers transformation (Eq. (2.45) in Ref. Giusto:2007fx , or Eq. (32) in Ref. Suzuki:2024eoz for precise expression), and Harrison transformation (Eq. (119) in Ref. Bouchareb:2007ax ), where the former adds the angular momenta to the five-dimensional asymptotically flat vacuum solution and the latter the electric charge.
@@ III Boundary conditions
In order that the solution (4) and (II) describes a capped black hole of physical interest, we must impose the appropriate boundary conditions at
@@ III.1 Asymptotic infinity ($x\to y\to-1$)
From the asymptotic behavior of the gauge potential (II), one can determine the electric charge
@@ III.2 Rotational axes ($y=-1,x=\pm 1$)
    To specifically ensure a topology of $S^{3}$, we must set $n=\pm 1$.
@@
    The condition for the absence of conical singularities: Since on this axis, the Killing vector $\partial_{\phi_{+}}:=\partial_{\phi}+n\partial_{\psi}$ vanishes under the two conditions previously mentioned, the inner axis does not exhibit a conical singularity if
@@ III.3 The center $(x,y)=(1,-1)$
where $t^{\prime}:=\sqrt{(1-\gamma)(1+\nu)}f_{1}f_{2}t/(c^{2}(1+\nu)f_{2}^{2}-s^{2}(1-\gamma)f_{1}^{2})$ and $(\psi_{+},\phi_{+})=(\psi-n\phi,\phi)$. Since the inequalities (19), (45) and (47) guarantees ${\mathcal{C}}_{1}>0$, the spacetime metric becomes the Minkowski metric. Therefore, the point $(x,y)=(1,-1)$ turns out to be a regular center of the Minkowski spacetime.
@@ III.4 Horizon
The spatial cross section of the horizon has topology $S^{3}$ under the boundary condition (34) with $n=\pm 1$. The area of the horizon cross section at $y=-1/\nu$ is given in terms of the surface gravity
@@
where $\omega_{\psi}^{+}=\omega_{\psi}-n\omega_{\phi}$ , $\omega_{\psi}^{-}=\omega_{\psi}$ , and $(\psi_{\pm},\phi_{\pm})$ are defined in Eq. (41). $R_{\pm}:=\sqrt{g_{\psi\psi}}|_{(x,y)=(\pm 1,-1/\nu)}$ are given by
@@ III.5 Absence of curvature singularities and CTCs
Figure 1: Profile of $H(x,y)$ for $(x,y)\in[-1,1]\times[-1/\nu,-1]$ and $(\sigma,\tilde{\nu},\tilde{\gamma})=(0.8,1.2,0.239)$ with $n=1$.
@@
Figure 2: Profile of ${\cal D}(x,y)$ for $(x,y)\in[-1,1]\times[-1/\nu,-1]$ and $(\sigma,\tilde{\nu},\tilde{\gamma})=(0.8,1.2,0.239)$ with $n=1$.
@@ III.6 Finding parameters for capped black holes
We are now aiming to find a capped black hole solution that does not have any curvature singularities, orbifold singularities, conical singularities, or Dirac-Misner string singularities both on and outside the horizon, additionally, they are free from closed timelike curves. This can be achieved if we can find a set of parameters $(\ell,\nu,\gamma,a,b,\beta,\alpha)$ that meets three constraints (33), (34) and (35). As outlined in our previous works Suzuki:2023nqf ; Suzuki:2024phv , such black holes are distinguished by having a compact inner rotational axis with a two-dimensional disk topology, referred to as a “bubble,” which lies adjacent to the horizon. The existence of this extra bubble implies that the intersection of the DOC and timeslice $\Sigma$ has the nontrivial topology of ${\rm DOC}\cap\Sigma=[{\mathbb{R}}^{4}\#{\mathbb{C}}{\mathbb{P}}^{2}]\setminus{\mathbb{B}}^{4}$, differing it from the Cvetič-Youm black hole with the trivial topology of ${\rm DOC}\cap\Sigma={\mathbb{R}}^{4}\setminus{\mathbb{B}}^{4}$ Cvetic:1996xz . We analyze the cases of $n=1$ and $n=-1$ separately.
Capped black hole with $n=1$.
@@
$\sigma^{3}=\frac{\tilde{b}\tilde{\nu}(\tilde{b}\tilde{\gamma}\tilde{\nu}+\tilde{b}^{2}\left(\tilde{\gamma}^{2}-2\right)+(\tilde{b}^{2}-1)\tilde{\nu}^{2}+2)}{\tilde{\gamma}(\tilde{b}(\tilde{b}^{2}-2)\tilde{\gamma}\tilde{\nu}+(1-2\tilde{b}^{2})\tilde{\nu}^{2}+2(\tilde{b}^{2}-1))}\frac{f_{2}\tilde{g}_{1}}{f_{1}g_{12}}.$ (81)
@@
Capped black hole with $n=-1$.
For $n=-1$, one can eliminate $\Delta,f_{1}$ and $f_{2}$ in Eq. (35) using Eq. (34) with $n=-1$, which leads to the condition
$\frac{(d_{1}-(1+a)(1-\gamma)(1+\nu)c_{1})^{2}}{4(1-\gamma)^{3}(1-\nu)^{2}(1+\nu)(1+a)^{2}}=1.$ (83)
In terms of $\tilde{\nu}$ and $\tilde{\gamma}$ defined in Eq. (71), this is solved with respect to $a$ as
$a=a_{\pm}:=\frac{\tilde{\gamma}^{2}\left(\tilde{\nu}^{2}-2\right)\left(\left(2b^{2}+b-1\right)\tilde{\nu}^{2}+2\right)+b^{2}\tilde{\nu}^{2}\left(\tilde{\nu}^{2}-2\right)^{2}+b(b+1)\tilde{\gamma}^{4}\tilde{\nu}^{2}\mp 2\tilde{\gamma}^{3}\tilde{\nu}\left(\tilde{\nu}^{2}-2\right)}{\tilde{\gamma}^{2}\tilde{\nu}\left((b+1)\tilde{\gamma}^{2}\tilde{\nu}+b\tilde{\nu}\left(\tilde{\nu}^{2}-2\right)\pm 2\tilde{\gamma}\left(\tilde{\nu}^{2}-2\right)\right)}.$ (84)
Because $a_{-}$ is derived from $a_{+}$ by the transformation $\tilde{\gamma}\to-\tilde{\gamma}$, it is sufficient to choose $a=a_{+}$ without loss of generality, by allowing $\tilde{\gamma}$ to take negative values as in the $n=1$ case. Substituting $a=a_{+}$ into the condition (45), one has
$d_{1}=-\frac{3(1-\tilde{b}^{2})\tilde{\gamma}^{2}\left(\tilde{\nu}^{2}-2\right)^{2}}{(\tilde{b}+2)^{2}}<0\quad\Longleftrightarrow\quad-1<\tilde{b}<1,$ (85)
where $\tilde{b}$ is defined by
$b=-\frac{\tilde{\gamma}\left(\tilde{\gamma}\tilde{\nu}+\tilde{b}(2-\tilde{\nu}^{2})\right)}{\tilde{\nu}\left(\tilde{\gamma}^{2}+\tilde{\nu}^{2}-2\right)}.$ (86)
Note that $\tilde{b}$ is defined differently from Eq. (77) in the $n=1$ case to ensure $\tilde{b}$ maintains the same parameter range With Eq. (86), $a=a_{+}$ in Eq. (84) and Eq. (33) reduce to
$a=\frac{(\tilde{b}^{2}-1)(\tilde{\nu}^{2}-2)}{(\tilde{b}+2)\tilde{\gamma}\tilde{\nu}}-1,$ (87)
$\sigma^{3}=-\frac{\tilde{\nu}\left(\left(\tilde{b}+2\right)\tilde{\gamma}\tilde{\nu}+\left(2\tilde{b}+1\right)\tilde{\gamma}^{2}-\left(\tilde{b}^{2}-1\right)\left(\tilde{\nu}^{2}-2\right)\right)}{\tilde{\gamma}\left(\left(2\tilde{b}^{2}+2\tilde{b}-1\right)\tilde{\gamma}\tilde{\nu}+\left(-\tilde{b}^{3}+2\tilde{b}+2\right)\tilde{\nu}^{2}+2\tilde{b}\left(\tilde{b}^{2}-1\right)\right)}\frac{\tilde{g}_{1}f_{2}}{f_{1}g_{12}},$ (88)
where we replaced $\alpha$ with $\sigma$ defined in Eq. (80). Finally, with Eqs. (86) and (87), Eq. (34) with $n=-1$ becomes
$\Delta\left[2(\tilde{b}+2)\tilde{\gamma}\tilde{\nu}+(\tilde{b}^{2}-1)\left(2-\tilde{\nu}^{2}\right))\right]-3f_{1}f_{2}\tilde{\gamma}\tilde{\nu}=0.$ (89)
Therefore, to obtain a physical solution, one must solve Eqs. (78) and (81) for $n=1$ or Eqs (88) and (89) for $n=-1$ under the conditions (19) and (46) for the parameters $(\tilde{\gamma},\tilde{\nu},\tilde{b},\sigma,\beta)$ within the specified range
@@
For $n=1$, since Eqs. (78) and (81) yield quartic and sixth-order equations of $\beta$, respectively, finding the analytic solution with respect to $\beta$ is not easy. Therefore, we resort to exploring the solutions numerically. The same is true for $n=-1$ with Eqs. (88) and (89). Since $\beta$ and $\tilde{b}$ are represented as the functions of $\sigma,\tilde{\nu}$ and $\tilde{\gamma}$ through Eqs. (78) and (81) for $n=1$ and through Eqs. (88) and (89) for $n=-1$, the physical solution is uniquely determined by a given $(\sigma,\tilde{\nu},\tilde{\gamma})$ in the range (90). In Fig. 3, we illustrate all possible solution for $n=1$, $\sigma=0.8,\ \tilde{\nu}=1.2$ and $0<|\tilde{\gamma}|<\sqrt{2-\tilde{\nu}^{2}}$. Moreover, we find that the metric and gauge potential is invariant under the transformation $\tilde{\gamma}\to-\tilde{\gamma}$, despite $\beta,\tilde{b}$ themselves not being invariant [^1] The metric function $H(x,y),F(x,y),J(x,y)$ are modified by the transformation $\tilde{\gamma}\to-\tilde{\gamma}$ but only upto a common factor $H(x,y)\to CH(x,y),\ F(x,y)\to CF(x,y),\ J(x,y)\to CJ(x,y)$. . As in the black ring case Suzuki:2024vzq , this might imply the possible extension from the $\tilde{\gamma}=0$ phase into the parameter region $\gamma\ (=1-\tilde{\gamma}^{2})>1$ where $\tilde{\gamma}$ becomes pure imaginary and $\beta,\tilde{b}$ become complex. In this paper, however, we restrict ourselves for the phases with all the parameters real.
It is worth mentioning the relationship between the two solutions for $n=1$ and $n=-1$. Since in a vacuum case, the reflection $x^{i}\to-x^{i}$ for either of the Killing coordinates $x^{i}=(t,\psi,\phi)$ transforms a solution into another with the angular momenta signs flipped, the transformation do not alter the physical or geometrical properties of the solutions. However, this is not the case in the five-dimensional minimal supergravity since the Maxwell-Chern-Simons equation (3) is not invariant under the reflection due to the presence of the Chern-Simons term. For example, let us consider the reflection $t\to-t$ for the $t$-component of Eq. (3):
$\begin{split}\partial_{x}(\sqrt{-g}F^{xt})+\partial_{y}(\sqrt{-g}F^{yt})+\frac{1}{4\sqrt{3}}(F_{\psi x}F_{\phi y}-F_{\phi x}F_{\psi y})=0.\end{split}$ (91)
This transformation flips the sign of first two terms, while keeping the last term unchanged, as it results in $F^{xt}\to-F^{xt},F^{yt}\to-F^{yt},F_{\psi x}\to F_{\psi x},F_{\phi x}\to F_{\phi x}$. Similarly, the other components of Eq. (3) are not invariant under $t\to-t$ either. Instead, one can observe that the following combination of the two reflections
$(t,\psi)\to(-t,-\psi)$ (92)
maintains Eq. (3) invariant. Since $g_{\psi\phi}$ and $A_{t}$ flip the signs, this converts the solution for $n=\pm 1$ and $\sigma$ into the solution for $n_{\rm new}=\mp 1$ and $\sigma_{\rm new}=-\sigma$, where the electric charge also flips, $Q\to Q^{\rm new}=-Q$. The mass and angular momenta transform as follows:
$(M,J_{\psi},J_{\phi})\to(M^{\rm new},J_{\psi}^{\rm new},J_{\phi}^{\rm new})=(M,J_{\psi},-J_{\phi})$ (93)
because $g_{t\phi}$ flips the sign while $g_{t\psi}$ remains unchanged. Therefore, the $n=1$ and $n=-1$ phases physically represent the same phase through the coordinate transformation (92). As previously mentioned, while all parameters remain real, both $n=1$ and $n=-1$ phases are incomplete and lack the region for $\gamma>1$. Nevertheless, it turns out that the transformation (92) maps the phase for $n=\pm 1$ with real parameters to the phase for $n=\mp 1$ in the region including the missing $\gamma>1$ phase (Fig 5). Hence, in the following section, we will focus on the $n=1$ phase but also use the solution for $n=-1$ to compensate the phase space for $\gamma>1$.
@@
Figure 3: All solutions for $n=1$, $\sigma=0.8$ and $\tilde{\nu}=1.2$ shown in the $(\tilde{\gamma},\beta)$ and $(\tilde{\gamma},\tilde{b})$ plains. $\beta$ diverges at $\tilde{\gamma}_{\infty}=-0.195...$ and the spacetime seems to be singular there. But the phases are smoothly continued at $\tilde{\gamma}=\tilde{\gamma}_{\infty}$ if we use the other parameter $\beta^{\prime}=\beta^{-1}$ around there. Despite $\beta=\beta(\tilde{\gamma})$ and $\tilde{b}=\tilde{b}(\tilde{\gamma})$ do not exhibit the mirror symmetry with respect to $\tilde{\gamma}$, the metric and gauge functions coincide for $\tilde{\gamma}=C>0$ (blue curve) and $\tilde{\gamma}=-C$ (red curve).